# gLSTM: Mitigating Over-Squashing by Increasing Storage Capacity

### v1

@@ 6 Conclusion
Many avenues remain open. Whereas the sensitivity issue of over-squashing has a mathematical basis via the node Jacobian, to our knowledge, the capacity issue does not. Theoretically quantifying this capacity could afford similar directions to those explored via sensitivity, establishing provable links to topology and model properties. With regards to the gLSTM architecture, we translate to a graph setting the gating and associative memory of xLSTM but do not retain the efficiency and parallel training, leaving open future work on more efficient MPNN architectures. Furthermore, another potential avenue would be to apply our findings to prevention of issues of over-mixing and representational collapse Barbero et al. (2024; 2025) in Transformer architectures.
@@
For easiest reproduction of our results, please consult the readme in the code repository provided in Appendix B. The repository includes all necessary information to run the experiments: in particular, configs containing the hyperparameters used (also reported in Section B.3) and code for all plots used in the paper.
@@ Acknowledgments
Appendix A xLSTM Update Equations
@@
Appendix B Additional Experiments
@@ B.3 Hyperparameters
B.4 Oversmoothing and Long Range Dependencies
@@
B.5 Additional NAR Classification Results
@@
B.5.1 Additional Sensitivity Metric Results
@@
B.6 Neighbor Associative Recall Regression Results
@@
Since the value vectors lack the sparsity of NAR, this appears to be a “harder” task in the sense that it is more taxing on memory capacity. This means that some of the over-squashing trends are more defined, particularly trends in sensitivity-based measures - see Section B.6.1. However, our experiments suggest that the regression target means that NARR becomes too hard for vector-memory MPNNs to effectively solve, visible in Figures 18 and 17.
@@
B.6.1 Relationship to Over-Squashing Sensitivity Metrics
@@
We hypothesize that the strong impact of the K-hop aggregation on the model’s ability to selectively recall - particularly visible for NARR - may partially explain the dramatic performance decrease when ablating this aggregation, discussed in Section B.2. We note that, while gLSTM consistently demonstrates superior performance to GCN, the improved performance is most striking when additionally using K-hop aggregation; it appears that the inductive bias introduced by the K-hop aggregation is particularly suited to the selective recall required by this task.

### v2

@@ 6 Conclusion
Many avenues remain open. Whereas the sensitivity issue of over-squashing has a mathematical basis via the node Jacobian, to our knowledge, the capacity issue does not. Theoretically quantifying this capacity could afford similar directions to those explored via sensitivity, establishing links to topology and model properties. With regards to architecture, we translate to a graph setting the gating and associative memory of xLSTM but do not retain the efficiency and parallel training, leaving open future work on more efficient MPNNs: we highlight that the recent work of Pöppel et al. (2025) achieves these efficiency gains in the specific case of directed acyclic graphs. Another potential avenue would be to apply our findings to prevention of issues of over-mixing and representational collapse (Barbero et al., 2024; 2025) in Transformer architectures.
@@
For easiest reproduction of our results, please consult the readme in the code repository provided in Appendix C. The repository includes all necessary information to run the experiments: in particular, configs containing the hyperparameters used (also reported in Section C.3) and code for all plots used in the paper.
@@ Acknowledgments
Appendix A Relationship to Computational (and Topological) Bottlenecks
Arnaiz-Rodriguez & Errica (2025) identify a separation in the over-squashing literature between the initial work of Alon & Yahav (2021) and the later work of Topping et al. (2022); Di Giovanni et al. (2023a). They point out that the original over-squashing definition of (Alon & Yahav, 2021) was associated with the computational tree, and Topping et al. (2022) later connected over-squashing to the existence of topological bottlenecks. They suggest that these definitions constitute different problems, and that – as a community – we should discard the term “over-squashing” and separate it into (at least) two separate terms: (1) computational tree bottlenecks and (2) topological bottlenecks (in the underlying graph).
This makes a valuable point: topological bottlenecks may or may not be involved when structural issues exist with the computational tree. The authors fairly point out that the umbrella term “over-squashing” has sometimes hidden some of the complexity of the problem, and the field may benefit from work identifying explicitly whether it is dealing with topological bottlenecks (as is the case with e.g. rewiring) or bottlenecks of the computational tree (e.g. adaptive message passing). We’re unsure that framing these as separate is helpful, since topological bottlenecks must be mediated through the computational tree in order to impact message passing, but the point stands that some methods to combat over-squashing specifically target topological bottlenecks (and as such any changes to the computational graph are implicit), and some methods target general computational bottlenecks, and that this is not always clear.
Despite the complexity, we believe the term “over-squashing” still has utility as an umbrella term that describes issues that arise from the presence of bottlenecks and depth in the computational tree. The distinction made by Arnaiz-Rodriguez & Errica (2025) clarifies that sometimes these issues arise due to topological bottlenecks in the underlying graph, and sometimes they do not. The relevance of our work is then in exploring how this structure manifests as performance issues, and the main body of our paper argues that this is due to separate issues of capacity and sensitivity.
Precisely how issues of capacity and sensitivity relate to the structure of the computational tree and the presence of bottlenecks is complex, although we take initial steps to clarify this in our paper. It would not be correct to suggest for example that capacity issues correspond to computational bottlenecks and low sensitivity to topological bottlenecks. In particular:
  • Our RingTransfer discussion of Section 3.1 highlights that computational bottlenecks can exist without capacity issues.
  • Arnaiz-Rodriguez & Errica (2025) highlight that low sensitivity does not necessarily imply the existence of computational tree bottlenecks (or topological bottlenecks).
  • On the other hand, both topological bottlenecks and general computational bottlenecks can cause low sensitivity. This is clear since the Equation 3 defines an upper bound on sensitivity in terms of the computational tree (defined by successive powers of the message-passing matrix), and we refer the reader to the work of Topping et al. (2022) for proofs in the case of specifically topological bottlenecks.
Additionally, while the discussion in this Appendix has so far focused on the computational tree, we highlight that the causes of the failure modes discussed in our work extend beyond structural issues: issues of capacity also depend at least on (1) information content of the task and (2) storage capacity of the model, as explored in Sections 3.1 and 3.2 respectively, and issues of sensitivity are highly dependent on the model dynamics, as explored in e.g. Arroyo et al. (2025); Heilig et al. (2025); Gravina et al. (2023; 2025).
Appendix B xLSTM Update Equations
@@ Appendix A Relationship to Computational (and Topological) Bottlenecks
Appendix C Additional Experiments
@@ C.3 Hyperparameters
C.4 Oversmoothing and Long Range Dependencies
@@
C.5 Additional NAR Classification Results
@@
C.5.1 Additional Sensitivity Metric Results
@@
C.6 Neighbor Associative Recall Regression Results
@@
Since the value vectors lack the sparsity of NAR, this appears to be a “harder” task in the sense that it is more taxing on memory capacity. This means that some of the over-squashing trends are more defined, particularly trends in sensitivity-based measures - see Section C.6.1. However, our experiments suggest that the regression target means that NARR becomes too hard for vector-memory MPNNs to effectively solve, visible in Figures 18 and 17.
@@
C.6.1 Relationship to Over-Squashing Sensitivity Metrics
@@
We hypothesize that the strong impact of the K-hop aggregation on the model’s ability to selectively recall - particularly visible for NARR - may partially explain the dramatic performance decrease when ablating this aggregation, discussed in Section C.2. We note that, while gLSTM consistently demonstrates superior performance to GCN, the improved performance is most striking when additionally using K-hop aggregation; it appears that the inductive bias introduced by the K-hop aggregation is particularly suited to the selective recall required by this task.